Calculate the approximate pH of an aqueous solution of H2S , 0.05 M, and H3PO4, 0.21 M at 25°C. (Note: consider one decimal place for the pH answer.)

pKa H2S/HS- = 7.02, pKa HS-/S2- = 13.99

pKaH3PO4/ H2PO4- =2.15, pKaH2PO4-/HPO42-=7.15, pKa HPO42-/ PO43- =12.38
Answer #1

Tst ioni3aonuaion Kal a:55%)。 = +H30 + バ = 1.02x10 14 PR92= 13-99 peal = -log (Ka) 13-99 Ka 7-02 lo Ka 억 H3 o + inthe SolutiI m Kal = 2. IS 0 00108 At iLibiym 0.21一乂 2 0.001481 0.00 70 sx = x 2. 2.-0.00708 士 2. -0.00708 + 0.0าา43 2 lo